OpenStudy (anonymous):

The formula for perimeter is P=2L+2W. Is this correctly expressed in terms of L? L=P-2W-2

OpenStudy (anonymous):

no, The quantity P - 2w must be divided by 2. P = 2L + 2w -2w -2w P-2w =2L /2 /2 (P-2w)/2 =L